OpenStudy (anonymous):

express log base2 64- log base2 4 as a single logarithm simplify if possible

OpenStudy (luigi0210):

Do you know your log rules?

OpenStudy (anonymous):

not really

OpenStudy (luigi0210):

Such as \(\large log_{a}x-log_{a}y=log_{a} \frac{x}{y}\)

OpenStudy (anonymous):

so would you just divide the two logs i have?

OpenStudy (anonymous):

\[64 = 2^{6}\] Therefore, \[\log _{2}64 = 6\] Similarly, \[\log _{2}4 = 2\] Therefore, result is 4.

OpenStudy (anonymous):

thanks